# Alexandrov spaces are CS sets

### v1

@@ 2.5. Extremal subsets
We give a more geometrical description of the proof of Claim 4.2 without relying on [8], [5] (at least explicitly). Let $\xi\in\Sigma_{p}$ and $r_{i}\to 0$. Consider the pointed Gromov–Hausdorff convergence $({r_{i}}^{-1}X,p)\to(T_{p},o)$, where $(T_{p},o)$ is the tangent cone at $p$ based at its vertex. Then there is a canonical way of lifting $\xi$ to $p_{i}\in{r_{i}^{-1}}S(p,r_{i})$ so that the distance function $d(p_{i},\cdot)$ is uniformly regular on a punctured ball around $p_{i}$ of fixed radius (after rescaling) independent of $i$ (the uniform regularity here means that the norm of the gradient of $d(p_{i},\cdot)$ is uniformly bounded below). See [24, 1.2], [36, 3.2] (actually a similar argument is included in the proof of the stability theorem used above). This, together with the fibration and stability theorems, implies that $\Sigma_{p_{i}}$ is homeomorphic to the space of direction of $T_{p}$ at $\xi$ (cf. [20, 5.1]). The latter is nothing but the spherical suspension over $\Sigma_{\xi}$, the space of directions of $\Sigma_{p}$ at $\xi$. Therefore, $\xi$ is a manifold point of $\Sigma_{p}$ if and only if $\Sigma_{p_{i}}$ is a topological sphere.
This canonical lifting argument obviously shows the “only if” part of Claim 4.2. The “if” part follows from Proposition 4.4(1) below and the following observation. Suppose another sequence $q_{i}\in S(p,r_{i})\setminus\{p_{i}\}$ converges to $\xi$. Then there exist three directions at $q_{i}$ making obtuse angles. This follows from the facts that $d(p_{i},\cdot)$ is uniformly regular at $q_{i}$ and that the two directions at $q_{i}$ to $p_{i}$ and $p$ are almost orthogonal, where the latter has an almost antipode. Then Proposition 4.4(1) shows that $\Sigma_{q_{i}}$ must be a topological sphere. Therefore, $p_{i}$ is the only candidate having a nonspherical space of directions among all sequences converging to $\xi$. This shows the “if” part of Claim 4.2.
@@
The following is a sphere theorem for an Alexandrov homology manifold of positive curvature and its link. The first half generalizes the double suspension theorem for an Alexandrov homology sphere in [35, p.750].
Proposition 4.4.
Let $\Sigma$ be an Alexandrov homology manifold with curvature $\geq 1$.
@@
    If $\Sigma$ contains a triple of points with pairwise distances $>\pi/2$, then it is homeomorphic to a sphere.
@@
    If $\Sigma$ contains a quadruple of points with pairwise distances $>\pi/2$, then every space of directions is homeomorphic to a sphere.
@@
We will follow the same strategy as in [17, Theorem C], which generalized Perelman’s diameter suspension theorem [25, 4.5] to the multiple suspension version in terms of the packing radius. In their terminology, our first/second assumption means that $\Sigma$ has the third/fourth packing radius $>\pi/4$, respectively.
(1) Let $(\xi_{1},\xi_{2},\xi_{3})$ be the triple of points in $\Sigma$ with pairwise distances $>\pi/2$. Then a standard triangle comparison argument shows that
  1. (i)
    there exists a unique point $\eta_{1}$ farthest from $\xi_{1}$;
  2. (ii)
    $\xi_{2},\xi_{3}$ are contained in the $\pi/2$-neighborhood of $\eta_{1}$; in particular, they make an obtuse angle at $\eta_{1}$;
  3. (iii)
    the distance function from $\eta_{1}$ has no critical points on $\Sigma\setminus\{\eta_{1},\zeta_{1}\}$, where $\zeta_{1}$ is the unique point farthest from $\eta_{1}$.
For the proofs of (ii) and (iii), see [17, 2.1, 2.8]. In particular, (iii) together with Perelman’s fibration theorem implies that $\Sigma$ is homeomorphic to a suspension with poles $\eta_{1},\zeta_{1}$. Moreover, (ii) shows that the pole $\eta_{1}$ is a manifold point as it is not an extremal point. This shows that $\Sigma$ is a topological sphere.
(2) Suppose there is a quadruple of points in $\Sigma$ with pairwise distances $>\pi/2$ but also there is a point of $\Sigma$ at which the space of directions is not homeomorphic to a sphere. By Theorem 4.1, the set $E$ of such singular points with nonspherical links is a $1$-dimensional extremal subset of $\Sigma$. Note that the proof of Theorem 4.1 shows that $E$ has no isolated points.
In fact any component $F$ of $E$ is a primitive extremal subset homeomorphic to a circle. Indeed, since $F$ is a compact MCS space, it is topologically a finite graph. If $F$ has a topological singularity $\xi$, i.e., a branching or boundary point, it must be an extremal point. However, since we have the quadruple, at least two of them are inside or outside the $\pi/2$-neighborhood of $\xi$. Then triangle comparison shows that $\xi$ is not an extremal point, a contradiction.
Similarly, for any $\xi\in F$, two of the quadruple lie inside the $\pi/2$-neighborhood of $\xi$ and the other two lie outside: otherwise there are three directions at $\xi$ making obtuse angles, which contradicts the first half. On the other hand, it is known that every point of $\Sigma$ is contained in the $\pi/2$-neighborhood of $F$ ([29, 1.4.1]). Therefore, if we go around the circle $F$, we will get a contradiction. ∎
@@
Both of the above statements are optimal in the sense that the numbers of points in the assumptions cannot be reduced. Indeed, the single and double suspensions of a homology sphere as in Example 2.22 give counterexamples to (1) and (2), respectively. On the other hand, if we assume $\Sigma$ to be a topological manifold from the beginning, then two points are enough to get the conclusion of (1), as in the original sphere theorem of Grove–Shiohama [16] (cf. [23, 8.2]).
The following is an advanced problem related to Question 3.4. An iterated space of directions refers to a space of directions of a space of directions of … of a space of directions.
Question 4.6.
Suppose an Alexandrov space contains no proper extremal subsets. Is any iterated space of directions homeomorphic to a sphere?
It is known that the above conclusion holds for any noncollapsing limit of Riemannian manifolds with sectional curvature bounded below, [19, 1.4] (though there may be proper extremal subsets). Yet another sufficient condition that implies the above conclusion is that every space of directions has radius greater than $\pi/2$, i.e., for any direction there is another direction making an obtuse angle with it, [30, §3], [31, §5.2] (cf. [15], [29, §1]; in this case there are no proper extremal subsets). See also [23, 1.3. 1.4, 1.5] for the corresponding results in the context of geodesically complete spaces with curvature bounded above.
@@
In this appendix we give an example of a primitive extremal subset of codimension $2$ whose induced intrinsic metric is not an Alexandrov metric with curvature bounded below. It is irrelevant to the main theorem, but like the main theorem, it cannot be found in the literature. This is why we include it here.
Let us briefly review the history. One of the long-standing problems in Alexandrov geometry is the boundary conjecture — the boundary of an Alexandrov space equipped with the induced intrinsic metric would have the same lower curvature bound as the ambient space. This is known for a convex hypersurface in a Riemannian manifold ([2]), but is still open even for a convex hypersurface in a smoothable Alexandrov space, i.e., a noncollapsing limit of Riemannian manifolds ([22]). Recall that the boundary is a codimension $1$ extremal subset.
As a generalization of the boundary conjecture, Perelman–Petrunin [29, 6.1] asked whether the induced intrinsic metric of a primitive extremal subset admits a lower curvature bound or not. In [30, §1 A counterexample], Petrunin constructed an example of a primitive extremal subset of codimension $3$ (and thus $\geq 3$) that is not an Alexandrov space (cf. [27, 4.1]). However, other than the boundary case, the codimension $2$ case remained. In [31, 9.1.2], Petrunin conjectured that a primitive extremal subset of codimension $2$ might be an Alexandrov space.
The following simple example provides a negative answer to this conjecture.
@@
Consider a convex subset $M$ of $\mathbb{R}^{4}$ defined by
$M:=\left\{(x,y,z,w)\in\mathbb{R}^{4}\mid x\geq 0,\ y\geq 0,\ z\geq 0\right\}.$
Suppose $\mathbb{Z}$ acts on $M$ by the cyclic permutation on the $(x,y,z)$-coordinate and by the translation on the $w$-coordinate. More precisely, it is generated by
@@
The quotient space $X:=M/\mathbb{Z}$ is an Alexandrov space of nonnegative curvature.
The $xw$-, $yw$-, and $zw$-planes are primitive extremal subsets of $M$. In particular, their union is an extremal subset of $M$, but not primitive. Let us denote by $E$ its projection into $X$, which is an extremal subset of codimension $2$ ([29, 4.1]). Then $E$ is primitive because it can no longer be divided into each plane due to the group action. However, $E$ cannot admit any Alexandrov metric with curvature bounded below as its intrinsic geodesics branch.
An extremal subset is called minimal if it contains no proper extremal subset. Clearly a minimal extremal subset is primitive. Note that any minimal extremal subset must be a topological manifold because the set of topological singularities, if nonempty, is a proper extremal subset ([29, 3.8]). In particular, the above counterexample is not minimal; indeed it contains the projection of the $w$-axis.
@@
Petrunin’s counterexample in [30, §1] is also not minimal: the vertex of the cone is an extremal point. However, it seems possible to slightly modify it so that the vertex becomes a nonextremal point.

### v2

@@ 2.6. Double suspension
More generally, Theorem 4.1 holds for any Alexandrov homology manifold $X$: the conclusion in this setting is that $E$ is an extremal subset of dimension $\leq 1$ (in particular, $E$ contains the isolated nonmanifold points of $X$). Let us trace the above proof, checking for required modifications. The dimension $\leq 4$ case follows from Properties (1)–(4), so we may assume dimension $\geq 5$ as before. Step 1 of the proof of Theorem 4.1 works exactly the same way. Step 2 requires the following minor modifications. In the proof of Claim 4.3, we used the assumption that $X$ is a topological manifold to ensure that $\Sigma_{x}$ is simply-connected (Property (3)), which implied that $\Sigma_{x}$ is a topological sphere if and only if it is a topological manifold. However, in the proof of Claim 4.3, $\Sigma_{x}$ is a topological suspension over $\Sigma$, where $\Sigma$ is connected since $\partial B(p,r)$ is a homology manifold. Therefore $\Sigma_{x}$ is automatically simply-connected. Another difference is that, in the last paragraph, Claim 4.3 may not imply that $E$ has no isolated points. We have to check that if $E$ has an isolated point $p$, then it is an extremal point. Indeed, Claim 4.3 implies that $\Sigma_{p}$ is a topological manifold, and thus it is not simply-connected (otherwise, it is a topological sphere). By Property (3), $p$ is an extremal point.
Remark 4.5.
Here we give a more geometric interpretation of the “only if” part of Claim 4.3. Let $\xi\in\Sigma_{p}$ and $r_{i}\to 0$. Consider the pointed Gromov–Hausdorff convergence $({r_{i}}^{-1}X,p)\to(T_{p},o)$, where $(T_{p},o)$ is the tangent cone of $X$ at $p$ based at the vertex. There is a canonical way of lifting $\xi$ to $p_{i}\in{r_{i}^{-1}}X$ so that the distance function $d(p_{i},\cdot)$ is uniformly regular around $p_{i}$ independent of $i$, that is,
$|\nabla d(p_{i},\cdot)|>\varepsilon$
on $r_{i}^{-1}\bar{B}(p_{i},r_{i}\varepsilon)\setminus\{p_{i}\}$, where $\varepsilon>0$ is independent of $i$ (see [35] for the definition of the gradient). For the proof, see [42, 3.2], [28, 1.2] (actually a similar argument is included in the proof of the stability theorem used above). This, together with the fibration and stability theorems, implies that $\Sigma_{p_{i}}$ is homeomorphic to $r_{i}^{-1}\partial B(p_{i},r_{i}\varepsilon)$. By the fibration and stability theorems again, $r_{i}^{-1}\partial B(p_{i},r_{i}\varepsilon)$ is homeomorphic to $\partial B(\xi,\varepsilon)\subset T_{p}$, which is homeomorphic to the space of directions of $T_{p}$ at $\xi$, provided $\varepsilon$ is small enough (cf. [23, 5.1]). This space of direction is isometric to the spherical suspension $S(\Sigma_{\xi})$, where $\Sigma_{\xi}$ is the space of directions of $\Sigma_{p}$ at $\xi$. Therefore,
$\xi$ is a manifold point of $\Sigma_{p}$
$\iff$ $\Sigma_{\xi}$ is simply-connected
[…]
This shows the “only if” part of Claim 4.3. It might also be possible to prove the “if” part by modifying this lifting argument, but we do not pursue it here.
@@
As an application of Theorem 4.1, we obtain a new sphere theorem for the link of an Alexandrov homology manifold with positive curvature.
Corollary 4.6.
Let $\Sigma$ be an Alexandrov homology manifold with curvature $\geq 1$. If $\Sigma$ contains four points with pairwise distance $>\pi/2$, then every space of directions is homeomorphic to a sphere.
Here $\Sigma$ itself is also a topological sphere, in particular, a topological manifold; see Remark 4.10(2). Hence Corollary 4.6 is an immediate consequence of Theorem 4.1 and the $k=1$ case of the following general lemma. By convention, the empty set has dimension $-1$.
Lemma 4.7.
Let $\Sigma$ be an Alexandrov space with curvature $\geq 1$. Suppose $\Sigma$ contains a $k$-dimensional extremal subset $F$, where $-1\leq k\leq\dim\Sigma$, satisfying the additional conditions as in Lemma 4.2 when $k=-1,0$. Then $\Sigma$ does not contain $k+3$ points with pairwise distance $>\pi/2$.
Note that the additional conditions above do not matter when proving Corollary 4.6, because the extremal subset obtained from Theorem 4.1 has dimension $1$.
Proof of Lemma 4.7.
Let $C(\Sigma)$ be the Euclidean cone over $\Sigma$. Then the subcone $C(F)$ is a $(k+1)$-dimensional extremal subset (this follows, for example, Lemma 4.2). If $\Sigma$ contains $k+3$ points with pairwise distance $>\pi/2$, then there exists a regular map at the vertex of $C(\Sigma)$ to $\mathbb{R}^{k+2}$. This contradicts Remark 2.20. ∎
Remark 4.8.
Corollary 4.6 is optimal in the sense that the number of discrete points and the inequality for the distance in the assumption cannot be relaxed. Indeed, the double suspension of a homology sphere as in Example 2.26 contains four points with pairwise distance $\geq\pi/2$.
Remark 4.9.
In general, if an Alexandrov space $\Sigma$ with curvature $\geq 1$ contains $k+1$ points with pairwise distance $>\pi/2$, then $\Sigma$ is homeomorphic to a $k$-fold suspension ([17, Theorem C]). In the maximal case, i.e., $k=\dim\Sigma+1$, we obtain that $\Sigma$ is bi-Lipschitz homeomorphic to the standard sphere (cf. [29, 3.2], [30]). In the Riemannian case, the existence of such $\pi/2$-discrete points gives some restriction on the differentiable structure ([18]).
Remark 4.10.
Let us recall what can be said if the number of $\pi/2$-discrete points is smaller (for the larger case, see Corollary 5.5 and Remark 5.6). An Alexandrov homology manifold $\Sigma$ with curvature $\geq 1$ is a topological sphere if one of the following holds:
@@
    $\Sigma$ contains no nonmanifold points and has diameter $>\pi/2$; or
@@
    $\Sigma$ contains three points with pairwise distance $>\pi/2$.
The first is a generalization of the diameter sphere theorem of Grove–Shiohama [16]. This follows from the suspension theorem and the basic fact that a sphere is the only manifold that is topologically a suspension (cf. [26, 8.2]). The second follows from the first, Property (2), and the following observation: under the assumption of (2), $\Sigma$ contains no extremal points. Indeed, for any $x\in\Sigma$, two of the three $\pi/2$-discrete points must lie inside or outside of the $\pi/2$-neighborhood of $x$, and thus they make an obtuse angle at $x$. Note that, as with Remark 4.8, the above assumptions are optimal.
5. Polyhedral case
Finally we prove Theorem 1.4. We first introduce polyhedral spaces. See [3, Ch. 12] (cf. [5, §3.2]) for more details. However, unlike [3, 12.1], we will only assume local compactness for polyhedral spaces instead of global compactness.
Let $\kappa\in\mathbb{R}$. The model space of curvature $\kappa$ refers to a sphere, Euclidean space, or a hyperbolic space of constant curvature $\kappa$. A complete geodesic space $X$ is called a polyhedral space if it admits a triangulation (i.e., a homeomorphism to a locally finite simplicial complex) such that each simplex in $X$ is isometric to a simplex in the model space of curvature $\kappa$, for some fixed $\kappa$. The triangulation of a polyhedral space means the triangulation in this definition. The cases $\kappa=1,0,-1$ are called, respectively, a spherical, Euclidean, hyperbolic polyhedral space. The dimension of a polyhedral space means the maximal dimension of a simplex (we only consider finite-dimensional polyhedral spaces). A characterization of an Alexandrov metric on a polyhedral space can be found in [3, 12.5] (cf. [5, 4.2.14]).
We first recall the basic properties of general (i.e., not necessarily Alexandrov) polyhedral metrics. Let $X$ be a polyhedral space. For a simplex $\sigma$ in $X$, the link around $\sigma$, denoted by $L(\sigma)$, is the subcomplex of $X$ formed by all the simplexes $\sigma^{\prime}$ such that $\sigma$ and $\sigma^{\prime}$ are disjoint but they are faces of a common simplex in $X$. The link $L(\sigma)$ is identified with the set of normal directions to $\sigma$ at some interior point of $\sigma$. Since the latter carries the angle metric, this makes $L(\sigma)$ a spherical polyhedral space.
Let $p\in X$. The space of directions $\Sigma_{p}X$ of $X$ at $p$ is defined in the same way as in Alexandrov geometry: the set of directions equipped with the angle metric (in this section the notation emphasizes the ambient space $X$). Note that $\Sigma_{p}X$ is a spherical polyhedral space of lower dimension. More precisely, if $p$ is contained in the interior of a $k$-dimensional simplex $\sigma$, then we have
$\Sigma_{p}X=S^{k}(L(\sigma)),$ (5.1)
where $S^{k}(\cdot)$ denotes the $k$-fold spherical suspension. Clearly a small neighborhood of $p$ is isometric to that of the vertex of the $\kappa$-cone over $\Sigma_{p}X$ (see [3, §11.C] or [5, 10.2.2] for the definition of $\kappa$-cone).
By definition, $\Sigma$ is a $k$-th iterated space of directions of $X$ at $p$ if there exists a sequence $p=\xi_{1},\dots,\xi_{k}$ such that
$\Sigma=\Sigma_{\xi_{k}}\cdots\Sigma_{\xi_{1}}X,\quad\xi_{i}\in\Sigma_{\xi_{i-1}}\cdots\Sigma_{\xi_{1}}X$
for any $2\leq i\leq k$. We denote this simply by
$\Sigma=\Sigma^{k}X$
without specifying the sequence $\xi_{1},\dots,\xi_{k}$.
A polyhedral space $X$ is PL (piecewise linear) if the link around every simplex $\sigma$ in $X$ is homeomorphic to a sphere. Note that any PL polyhedral space is in particular a topological manifold. If a polyhedral space is PL, then every link and space of directions are again PL. Furthermore, the following holds.
Lemma 5.1.
Let $X$ be a polyhedral space. Then $X$ is PL if and only if every iterated space of directions is homeomorphic to a sphere.
Therefore, in the polyhedral case, Conjecture 1.3 is equivalent to Theorem 1.4.
@@ 5. Polyhedral case
The “if” part follows by taking iterated spaces of directions of both sides of Equation (5.1). The “only if” part is proved by induction on dimension. If $X$ is PL, Equation (5.1) shows that every space of directions is homeomorphic to a sphere. Since the link is also PL, the inductive assumption and Equation (5.1) show that every $k$-th iterated space of directions ($k\geq 2$) is homeomorphic to a sphere. ∎
Finally, we give a direct characterization of an extremal subset in a polyhedral Alexandrov space. Let $X$ be a polyhedral Alexandrov space. Then a subset $E$ of $X$ is extremal (in the sense of Definition 2.14) if and only if
  • $E$ is a subcomplex of $X$; and
  • for every simplex $\sigma$ in $E$, if $\xi\notin B(L_{E}(\sigma),\pi/2)$, then $L(\sigma)\subset\bar{B}(\xi,\pi/2)$.
Here $L_{E}(\sigma)$ denotes the link of $\sigma$ in the subcomplex $E$, which is a subcomplex of $L(\sigma)$. The proof is straightforward and left to the reader (we will not use it).
We now prove Theorem 1.4. We show the following more rigid result. Compare with Theorem 4.1 and Remark 4.4.
Theorem 5.2.
Let $X$ be a polyhedral Alexandrov homology $n$-manifold. For any integer $1\leq k\leq n$, let $E_{k}$ be the set of points $p\in X$ such that some $k$-th iterated space of directions at $p$ is not homeomorphic to $S^{n-k}$. If $E_{k}$ is nonempty, then it is an extremal subset of $X$ that is a subcomplex of $X$ only with $k$- or $(k-1)$-dimensional maximal simplexes.
By Properties (1) and (4) in Section 4, $E_{k}$ is empty for $k\geq n-2$, so we may assume $k\leq n-3$. Moreover, we have already proved in Theorem 4.1 and Remark 4.4 that $E_{1}$ is an extremal subset of dimension $\leq 1$. In the polyhedral case, we also see that $E_{1}$ is a subcomplex of $X$, since the space of directions is invariant in the interior of a simplex as in Equation (5.1). Therefore we may assume $k\geq 2$ (and hence $n\geq 5$).
We prove Theorem 5.2 by induction on dimension (more precisely, on $k$). Recall that, if $X$ is a polyhedral Alexandrov homology manifold, then $\Sigma_{p}X$ is again an Alexandrov polyhedral homology manifold with the homology of a sphere. The key observation is the following. Compare with Claim 4.3.
Claim 5.3.
Let $X$ be as in Theorem 5.2 and $p\in X$. Then we have
$\Sigma_{p}E_{k}(X)=E_{k-1}(\Sigma_{p}X)$
for any $2\leq k\leq n$, where $E_{k}(\cdot)$ denotes the singular set defined in Theorem 5.2. The equality remains valid even if either side is empty.
@@
Claim 4.3 showed that
$\Sigma_{p}E_{1}(X)=E_{0}(\Sigma_{p}X),$
where $E_{0}(\cdot)$ denotes the set of nonmanifold points. The following proof of Claim 5.3 also works for $k=1$, which is indeed a baby version of the proof of Claim 4.3. Theorem 5.2 also holds for $E_{0}$ (Property (2) in Section 4).
Proof of Claim 5.3.
The proof only uses the properties of polyhedral spaces and does not rely on the properties of Alexandrov spaces. Let $U$ be a neighborhood of $p$ that is isometric to a neighborhood of the vertex of the $\kappa$-cone over $\Sigma_{p}X$. Let $x\in U$ be a point in the direction $\xi\in\Sigma_{p}X$. Then we have
$\xi\in\Sigma_{p}E_{k}(X)\iff x\in E_{k}(X)\iff\Sigma^{k-1}\Sigma_{x}X\ncong S^{n-k},$ (5.2)
where $\cong$ denotes a homeomorphism.
Since $\Sigma_{x}X=S(\Sigma_{\xi}\Sigma_{p}X)$, we have the following two possibilities:
$\Sigma^{k-1}\Sigma_{x}X=\begin{cases}\hfil\Sigma^{k-2}\Sigma_{\xi}\Sigma_{p}X\\ S(\Sigma^{k-1}\Sigma_{\xi}\Sigma_{p}X)\end{cases}$
(consider the $k=2$ case and apply it repeatedly). In the first case, it is clear that $\Sigma^{k-1}\Sigma_{x}X$ is not homeomorphic to $S^{n-k}$ if and only if $\xi\in E_{k-1}(\Sigma_{p}X)$. In the second case, $\Sigma^{k-1}\Sigma_{x}X$ is not homeomorphic to $S^{n-k}$ if and only if $\Sigma^{k-1}\Sigma_{\xi}\Sigma_{p}X$ is not simply-connected, by Edwards’ theorem [9, §8] (or Property (3) in Section 4). The latter is equivalent to that $\Sigma^{k-2}\Sigma_{\xi}\Sigma_{p}X$ is not a manifold, i.e., $\xi\in E_{k-1}(\Sigma_{p}X)$. Therefore, in either case,
$\Sigma^{k-1}\Sigma_{x}X\ncong S^{n-k}\iff\xi\in E_{k-1}(\Sigma_{p}X).$ (5.3)
Combining (5.2) and (5.3), we obtain the conclusion. ∎
Proof of Theorem 5.2.
We prove this by induction on $k$. The base case, $k=1$, was already proved in Theorem 4.1 and Remark 4.4. Suppose $k\geq 2$.
We first show that $E_{k}(X)$ is a subcomplex (in particular, a closed subset of $X$). This follows from Claim 5.3. If $p$ is contained in the closure of $E_{k}(X)$, then Claim 5.3 implies that $E_{k-1}(\Sigma_{p}X)$ is nonempty. This means that $p\in E_{k}(X)$, i.e., $E_{k}(X)$ is closed. If the interior of a simplex $\sigma$ intersects $E_{k}(X)$, then clearly every point in the interior is contained in $E_{k}(X)$. Since $E_{k}(X)$ is closed, the boundary of $\sigma$ is contained in $E_{k}(X)$, i.e., $E_{k}(X)$ is a subcomplex.
By the inductive assumption and Claim 5.3, for any $p\in E_{k}(X)$, the space of directions $\Sigma_{p}E_{k}(X)$ is an extremal subset of $\Sigma_{p}X$ that is a subcomplex of $\Sigma_{p}X$ only with $(k-1)$- or $(k-2)$-dimensional maximal simplexes. This implies that $E_{k}(X)$ only have $k$- or $(k-1)$-dimensional maximal simplexes. Furthermore, by Lemma 4.2, $E_{k}(X)$ is an extremal subset of $X$. Note that the additional conditions in Lemma 4.2 are satisfied. Indeed, Claim 5.3 implies that $\Sigma_{p}E_{k}(X)$ is nonempty (since $k\geq 2$), and if $\Sigma_{p}E_{k}(X)$ consists of one element (when $k=2$), the same argument as in Step 1 of the proof of Claim 4.3 shows that $\operatorname{diam}\Sigma_{p}E_{k}(X)\leq\pi/2$. This completes the proof. ∎
The following corollary immediately follows from Theorem 5.2 and Lemma 4.7. Compare with Corollary 4.6.
Corollary 5.5.
Let $\Sigma$ be a polyhedral Alexandrov homology manifold with curvature $\geq 1$. Suppose $\Sigma$ contains $k+3$ points with pairwise distance $>\pi/2$. Then every $k$-th iterated space of directions is homeomorphic to a sphere.
Remark 5.6.
The above proof of Theorem 5.2 (except for the part about polyhedral structures) only uses general properties of Alexandrov spaces. Therefore the proof of Conjecture 1.3 in the general case reduces to the proof of Claim 5.3. Similarly, if we obtain Theorem 5.2 for nonpolyhedral spaces, then it immediately implies Corollary 5.5 in the nonpolyhedral case.
Finally, we remark that the absence of extremal subsets, as in Conjecture 1.3, is not inherited to the space of directions (if it were, the proof of Conjecture 1.3 would be straightforward by induction on dimension).
Example 5.7.
Let $P$ be a truncated cube obtained by cutting off all the corners of a solid cube by planes. Let $p\in P$ a midpoint of an edge of the original cube. Consider the double $D(P)$ of $P$, i.e., the gluing of two copies of $P$ along the boundary. Then $D(P)$ contains no proper extremal subsets. On the other hand, the space of directions $\Sigma_{p}D(P)$ is isometric to the double of the quadrant of the unit sphere, which contains a 00-dimensional extremal subset. A similar example can be obtained by taking the intersection of the solid cube in $\mathbb{R}^{3}$ with edge length $2$ and the concentric Euclidean ball of radius $\sqrt{2}$ and doubling it. In this case the resulting space has no proper extremal subsets even “locally” around the midpoint of the original edge.
@@
In this appendix, we give an example of a primitive extremal subset of codimension $2$ whose induced intrinsic metric is not an Alexandrov metric with a lower curvature bound. The construction is very simple, but to the best of the author’s knowledge, no such examples can be found in the literature.
Let us briefly review the history. One of the long-standing problems in Alexandrov geometry is the boundary conjecture — the boundary of an Alexandrov space equipped with the induced intrinsic metric would have the same lower curvature bound as the ambient space. This is true for a convex hypersurface in a Riemannian manifold ([2]), but is still open even for a convex hypersurface in a smoothable Alexandrov space, i.e., a noncollapsing limit of Riemannian manifolds with sectional curvature bounded below ([25, p.3871]). Recall that the boundary is a codimension $1$ extremal subset.
As a generalization of the boundary conjecture, Perelman–Petrunin [33, 6.1] asked whether the induced intrinsic metric of a primitive extremal subset admits a lower curvature bound or not. In [34, §1], Petrunin constructed an example of a primitive extremal subset of codimension $3$ (and thus $\geq 3$) that is not an Alexandrov space (cf. [31, 4.1]). However, other than the boundary case, the codimension $2$ case remained. In [35, 9.1.2], Petrunin conjectured that a primitive extremal subset of codimension $2$ might be an Alexandrov space.
The following example provides a negative answer to this conjecture.
@@
Let $X$ be a convex subset of $\mathbb{R}^{4}$ defined by
$X:=\left\{(x,y,z,w)\in\mathbb{R}^{4}\mid x\geq 0,\ y\geq 0,\ z\geq 0\right\}.$
Suppose $\mathbb{Z}$ acts isometrically on $X$ by the cyclic permutation on the $(x,y,z)$-coordinate and by the translation on the $w$-coordinate. More precisely, the action is generated by
@@
Then the quotient space $X/\mathbb{Z}$ is an Alexandrov space with nonnegative curvature. Let $\pi:X\to X/\mathbb{Z}$ be the projection.
The $xw$-, $yw$-, and $zw$-planes are primitive extremal subsets of $X$. In particular, their union $E$ is an extremal subset of $X$, but not primitive. The projection $\pi(E)$ is an extremal subset of codimension $2$ in $X/\mathbb{Z}$ ([33, 4.1]). Furthermore, $\pi(E)$ is primitive because it can no longer be divided into each plane, due to the effect of the group action. However, $\pi(E)$ does not admit any Alexandrov metric with a lower curvature bound as its intrinsic geodesics branch.
An extremal subset is called minimal if it contains no proper extremal subsets. Clearly a minimal extremal subset is primitive. Note that any minimal extremal subset must be a topological manifold, because the set of topological singularities, if nonempty, is a proper extremal subset. In particular, the above counterexample is not minimal: the projection of the $w$-axis is a proper extremal subset of $\pi(E)$.
@@
Petrunin’s counterexample in [34, §1] is also not minimal: the vertex of the cone is an extremal point. However, it is possible to slightly modify it so that the vertex becomes a nonextremal point.